1.    The diameter of holes for a cable harness is known to have a normal distribution with a standard deviation of 0.01 inch. A random sample of size 10 yields an average diameter of 1.5045 inch. 

a.    Find a 99% two-sided confidence interval on the mean hole diameter.


b.    Find a 90% two-sided confidence interval on the mean hole diameter.


c.     Explain the difference of the interval between the two confidence levels.



1
Expert's answer
2021-09-13T00:04:43-0400

"n=10 \\

M = 1.5045 \\

\sigma=0.01 \\

CI = (M - \frac{Z_c \times \sigma}{\sqrt{n}}, M + \frac{Z_c \times \sigma}{\sqrt{n}})"

a.

"Z_c = 2.576 \\

CI = (1.5045 - \frac{2.576 \times 0.01}{\sqrt{10}}, 1.5045 + \frac{2.576 \times 0.01}{\sqrt{10}}) \\

=(1.5045 - 0.008146, 1.5045 + 0.008146) \\

=(1.496354, 1.512746)"

b.

"Z_c = 1.645 \\

CI = (1.5045 - \frac{1.645 \times 0.01}{\sqrt{10}}, 1.5045 + \frac{1.645 \times 0.01}{\sqrt{10}}) \\

=(1.5045 - 0.005202, 1.5045 + 0.005202) \\

=(1.499298, 1.509702)"

c. A 99 percent confidence interval is wider than a 90 percent confidence interval because to be more confident that the true population value falls within the interval we will need to allow more potential values within the interval.